OpenStudy (anonymous):

Suppose that in one week, your company takes in $2,220 in sales at one of its cell phone stores. The price breakdown per phone is as follows: Cell phone model A4: $35 Smartphone Z20: $50 Suppose that a total of 51 phones were sold. this is the supposed problem with the method of solving in a graph

OpenStudy (anonymous):

Standard form: a+z=51; 35a+50z=2220; substitute/eliminate a = -1z+51 substitute/eliminate z = +29 Solution: a = 22; z = 29;